# Faster Cascades via Speculative Decoding

### v1

@@ A.5 Proof of Lemma 5
For a fixed prefix $x_{<t}$, we can write the deferral risk in (8) as:
@@
$~{}~{}~{}~{}~{}+\left(\hat{r}_{\rm\tt OPT}(x_{<t})-r^{*}(x_{<t})\right)\cdot\left(\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},p_{t})\right]-\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},q_{t})\right]-\max_{v}q_{t}(v)+\max_{v}p_{t}(v)\right)$
@@
$~{}~{}~{}~{}~{}+\left(\hat{r}_{\rm\tt OPT}(x_{<t})-r^{*}(x_{<t})\right)\cdot\left(\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},p_{t})\right]-1+\max_{v}p_{t}(v)\right)$
$~{}~{}~{}~{}~{}+\left(\hat{r}_{\rm\tt OPT}(x_{<t})-r^{*}(x_{<t})\right)\cdot\left(1-\max_{v}q_{t}(v)-\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},q_{t})\right]\right)$
@@
$~{}~{}~{}~{}~{}+\left|\hat{r}_{\rm\tt OPT}(x_{<t})-r^{*}(x_{<t})\right|\cdot\left|\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},p_{t})\right]-1+\max_{v}p_{t}(v)\right|$
$~{}~{}~{}~{}~{}+\left|\hat{r}_{\rm\tt OPT}(x_{<t})-r^{*}(x_{<t})\right|\cdot\left|1-\max_{v}q_{t}(v)-\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},q_{t})\right]\right|$
@@
$~{}~{}~{}~{}~{}+\underbrace{\left|\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},p_{t})\right]-1+\max_{v}p_{t}(v)\right|}_{\text{term}_{2}}+\underbrace{\left|1-\max_{v}q_{t}(v)-\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},q_{t})\right]\right|}_{\text{term}_{3}}$ (13)
@@
$=\left|\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[1\left(x_{t}\neq\operatorname{argmax}_{v}p_{t}(v)\right)\right]-1+\max_{v}p_{t}(v)\right|$
$=\left|\max_{v}p_{t}(v)-\sum_{x_{t}}\mathbb{P}(x_{t}|x_{<t})\cdot 1\left(x_{t}=\operatorname{argmax}_{v}p_{t}(v)\right)\right|$
Suppose $v^{*}\in\operatorname{argmax}_{v}p_{t}(v)$, then:
@@
Substituting (14)–(16) in (13) completes the proof.
@@
SpecDecoding [26] - $p(x)$ Speculative
Lossy SpecDecoding [44] - $\max\{\min\{p(x),\frac{q(x)}{1-\alpha}\},\frac{q(x)}{\beta}\}$ Speculative
BiLD* [24] $\mathbf{1}\big{(}\,D(q,p)>\alpha\big{)}$ $(1-\delta)\cdot q(x)+\delta\cdot p(x)$ Speculative
Cascade [Chow] [10] $\mathbf{1}\big{(}\max_{v}q(v)<1-\alpha\big{)}$ $(1-\delta)\cdot q(x)+\delta\cdot p(x)$ Sequential
Cascade [ChowLog] $\mathbf{1}\big{(}\,\textrm{entropy}(q)>\alpha\big{)}$ $(1-\delta)\cdot q(x)+\delta\cdot p(x)$ Sequential
Oracle [Diff] [22] $\mathbf{1}\big{(}\max_{v}q(v)<\max_{v}p(v)-\alpha\big{)}$ $(1-\delta)\cdot q(x)+\delta\cdot p(x)$ Oracle
Oracle [DiffLog] $\mathbf{1}\big{(}\,\textrm{entropy}(p)<\textrm{entropy}(q)-\alpha\big{)}$ $(1-\delta)\cdot q(x)+\delta\cdot p(x)$ Oracle
SpecCascade [Chow] $\mathbf{1}\big{(}\max_{v}q(v)<1-\alpha\big{)}$ $(1-\delta)\cdot q(x)+\delta\cdot p(x)$ Speculative
SpecCascade [ChowLog] $\mathbf{1}\big{(}\,\textrm{entropy}(q)>\alpha\big{)}$ $(1-\delta)\cdot q(x)+\delta\cdot p(x)$ Speculative
SpecCascade [Diff01] $\mathbf{1}\big{(}\max_{v}q(v)<\max_{v}p(v)-\alpha\big{)}$ $(1-\delta)\cdot q(x)+\delta\cdot p(x)$ Speculative
SpecCascade [DiffLog] $\mathbf{1}\big{(}\,\textrm{entropy}(p)<\textrm{entropy}(q)-\alpha\big{)}$ $(1-\delta)\cdot q(x)+\delta\cdot p(x)$ Speculative
SpecCascade [OPT01] $\mathbf{1}\big{(}\max_{v}q(v)<\max_{v}p(v)-\alpha\cdot D_{\textrm{TV}}(p,q)\big{)}$ $(1-\delta)\cdot q(x)+\delta\cdot p(x)$ Speculative
SpecCascade [OPTLog] $\mathbf{1}\big{(}\,\textrm{entropy}(p)<\textrm{entropy}(q)-\alpha\cdot D_{\textrm{TV}}(p,q)\big{)}$ $(1-\delta)\cdot q(x)+\delta\cdot p(x)$ Speculative
Table 3: Target distributions associated with different inference algorithms, where $\alpha$ is a free parameter and $\beta\geq 1-\alpha$ is a parameter dependent on $q,p$ and $\alpha$. The last column indicates whether the execution is sequential (Algorithm 2), via an oracle (Algorithm 3), or speculative (Algorithm 5) [26]. The third row presents a variant of the BiLD algorithm of [24], where $D(q,p)$ is a measure of discrepancy between $q$ and $p$; the original algorithm differs from [26] in the use of a deterministic speculative decoding procedure with a dynamic draft window (see §5).
@@ Appendix B Derivation of Chow’s rule
The minimizer of (17) is of the form:
$r^{*}(x_{<t})=1~{}~{}\iff~{}~{}\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},q_{t})\right]\,>\,\alpha.$ (18)
If $\ell=\ell_{\text{0-1}}$, one may employ a plug-in estimator to (18) by replacing the expected 0-1 loss over $q_{t}$ with $1-\max_{v}q_{t}(v)$, giving us $\hat{r}_{\rm\tt Chow}(x_{<t})$ in (2). If $\ell=\ell_{\log}$, one may replace the expected log loss over $q_{t}$ with the entropy of $q_{t}$, giving us:
$\hat{r}_{\rm\tt ChowLog}(x_{<t})=1~{}\iff~{}\textrm{entropy}\big{(}q(\cdot|x_{<t})\big{)}>\alpha,$ (19)
@@ Appendix C Optimal Deferral: Additional Discussion
We provide additional discussion for the optimal deferral rules derived in §3 and §4.3.
@@ C.1 Optimal sequential deferral when $\ell=\ell_{\log}$
$r^{*}(x_{<t})=1~{}~{}\iff~{}~{}\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},q_{t})\right]\,>\,\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},p_{t})\right]+\alpha\cdot D_{\textup{{TV}}}(p_{t},q_{t}).$
@@
$\textstyle\hat{r}_{\rm\tt DiffLog}(x_{<t})=1~{}~{}\iff~{}~{}\sum_{v}q_{t}(v)\cdot\log(q_{t}(v))\,<\,\sum_{v}p_{t}(v)\cdot\log(p_{t}(v))-\alpha.$ (20)
@@ C.2 Optimal speculative deferral when $\ell=\ell_{\log}$
$r^{*}(x_{<t})=1~{}~{}\iff~{}~{}\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},q_{t})\right]\,>\,\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},p_{t})\right]+\alpha\cdot D_{\textup{{TV}}}(p_{t},q_{t}).$
@@
$\textstyle\hat{r}_{\rm\tt OPTLog}(x_{<t})=1~{}\iff~{}\sum_{v}q_{t}(v)\cdot\log(q_{t}(v))\,<\,\sum_{v}p_{t}(v)\cdot\log(p_{t}(v))-\alpha\cdot D_{\textup{{TV}}}(p_{t},q_{t}).$ (21)
@@
$L_{\rm spec}(r_{\rm\tt OPT};x_{<t})-\min_{r}L_{\rm spec}(r;x_{<t})~{}\leq~{}\textstyle B_{q}\cdot\sum_{v\in\mathscr{V}}\big{|}\mathbb{P}(v|x_{<t})-q_{t}(v)\big{|}\,+\,B_{p}\cdot\sum_{v\in\mathscr{V}}\big{|}\mathbb{P}(v|x_{<t})-p_{t}(v)\big{|}.$
@@
Plugging these bounds into the equivalent of (13) in Lemma 5 for the log-loss completes the proof. ∎
@@ C.3 Optimal speculative deferral for greedy decoding
When $T\rightarrow 0$, running Algorithm 5 with $\tilde{r}_{\rm\tt OPT}$ as the deferral rule and $\tilde{q}_{t}$ as the drafter is equivalent to running it with $\hat{r}_{\rm\tt Diff}$ in (5) as the deferral rule and $\tilde{q}_{t}$ as the drafter.
@@
Note that under greedy inference, $\tilde{q}_{t}$ $\tilde{p}_{t}$ are one-hot encodings of $\operatorname{argmax}_{v}q_{t}(v)$ and $\operatorname{argmax}_{v}p_{t}(v)$ respectively. As a result,
$D_{\textrm{TV}}(\tilde{q}_{t},\tilde{p}_{t})=\mathbf{1}\left(\operatorname{argmax}_{v}q_{t}(v)\neq\operatorname{argmax}_{v}p_{t}(v)\right).$
@@
where $\delta_{\rm OPT}(q,p)~{}=~{}\mathbf{1}\left(\max_{v}q(v)\,<\,\max_{v}p(v)-\alpha\cdot\mathbf{1}\left(\operatorname{argmax}_{v}q(v)\neq\operatorname{argmax}_{v}p(v)\right)\right)$. When $\operatorname{argmax}_{v}q(v)=\operatorname{argmax}_{v}p(v)$, then $\tilde{q}=\tilde{p}$, and irrespective of the outcome of $\delta(q,p),$ we have that $\pi(v)=1$. When $\operatorname{argmax}_{v}q(v)\neq\operatorname{argmax}_{v}p(v)$, then
$\pi(v)=1-\delta_{\rm OPT}(q,p)=\mathbf{1}\left(\max_{v}q(v)\,\geq\,\max_{v}p(v)-\alpha\right)=1-\delta_{\rm\tt Diff}(q,p).$
@@
When $\operatorname{argmax}_{v}q(v)=\operatorname{argmax}_{v}p(v)$, $p_{\rm res}(v)=0$. When $\operatorname{argmax}_{v}q(v)\neq\operatorname{argmax}_{v}p(v)$,
@@
C.4 Equivalence between (7) and (8)
Since the prefix $x_{<t}$ is fixed in (7), the constrained optimization we seek to solve is of essentially of the following form:
$\min_{r\in\{0,1\}}(1-r)\cdot c_{0}+r\cdot c_{1}~{}~{}~{}\text{s.t.}~{}~{}~{}r\cdot c_{2}\leq B,$

### v2

@@ A.5 Proof of Lemma 5
For a fixed prefix $x_{<t}$, we can write the deferral risk in equation 8 as:
@@
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\left(\hat{r}_{\rm\tt OPT}(x_{<t})-r^{*}(x_{<t})\right)\cdot\left(\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},p_{t})\right]-\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},q_{t})\right]-\max_{v}q_{t}(v)+\max_{v}p_{t}(v)\right)$
@@
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\left(\hat{r}_{\rm\tt OPT}(x_{<t})-r^{*}(x_{<t})\right)\cdot\left(\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},p_{t})\right]-1+\max_{v}p_{t}(v)\right)$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\left(\hat{r}_{\rm\tt OPT}(x_{<t})-r^{*}(x_{<t})\right)\cdot\left(1-\max_{v}q_{t}(v)-\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},q_{t})\right]\right)$
@@
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\left|\hat{r}_{\rm\tt OPT}(x_{<t})-r^{*}(x_{<t})\right|\cdot\left|\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},p_{t})\right]-1+\max_{v}p_{t}(v)\right|$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\left|\hat{r}_{\rm\tt OPT}(x_{<t})-r^{*}(x_{<t})\right|\cdot\left|1-\max_{v}q_{t}(v)-\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},q_{t})\right]\right|$
@@
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\underbrace{\left|\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},p_{t})\right]-1+\max_{v}p_{t}(v)\right|}_{\text{term}_{2}}+\underbrace{\left|1-\max_{v}q_{t}(v)-\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},q_{t})\right]\right|}_{\text{term}_{3}}$ (18)
@@
$=\left|\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[1\left(x_{t}\neq\operatorname*{arg\,max}_{v}p_{t}(v)\right)\right]-1+\max_{v}p_{t}(v)\right|$
$=\left|\max_{v}p_{t}(v)-\sum_{x_{t}}\mathbb{P}(x_{t}|x_{<t})\cdot 1\left(x_{t}=\operatorname*{arg\,max}_{v}p_{t}(v)\right)\right|$
Suppose $v^{*}\in\operatorname*{arg\,max}_{v}p_{t}(v)$, then:
@@
Substituting equation 19–equation 21 in equation 18 completes the proof.
@@
SpecDecoding Leviathan et al. (2023) - $p(x)$ Speculative
Lossy SpecDecoding (Tran-Thien, 2023) - $\max\{\min\{p(x),\frac{q(x)}{1-\alpha}\},\frac{q(x)}{\beta}\}$ Speculative
BiLD* (Kim et al., 2023) $\bm{1}\big{(}\,D(q,p)>\alpha\big{)}$ $(1-\delta)\cdot q(x)+\delta\cdot p(x)$ Speculative
Cascade [Chow] (Chow, 1970) $\bm{1}\big{(}\max_{v}q(v)<1-\alpha\big{)}$ $(1-\delta)\cdot q(x)+\delta\cdot p(x)$ Sequential
Cascade [ChowLog] $\bm{1}\big{(}\,\textrm{entropy}(q)>\alpha\big{)}$ $(1-\delta)\cdot q(x)+\delta\cdot p(x)$ Sequential
Oracle [Diff] (Jitkrittum et al., 2023) $\bm{1}\big{(}\max_{v}q(v)<\max_{v}p(v)-\alpha\big{)}$ $(1-\delta)\cdot q(x)+\delta\cdot p(x)$ Oracle
Oracle [DiffLog] $\bm{1}\big{(}\,\textrm{entropy}(p)<\textrm{entropy}(q)-\alpha\big{)}$ $(1-\delta)\cdot q(x)+\delta\cdot p(x)$ Oracle
SpecCascade [Chow] $\bm{1}\big{(}\max_{v}q(v)<1-\alpha\big{)}$ $(1-\delta)\cdot q(x)+\delta\cdot p(x)$ Speculative
SpecCascade [ChowLog] $\bm{1}\big{(}\,\textrm{entropy}(q)>\alpha\big{)}$ $(1-\delta)\cdot q(x)+\delta\cdot p(x)$ Speculative
SpecCascade [Diff01] $\bm{1}\big{(}\max_{v}q(v)<\max_{v}p(v)-\alpha\big{)}$ $(1-\delta)\cdot q(x)+\delta\cdot p(x)$ Speculative
SpecCascade [DiffLog] $\bm{1}\big{(}\,\textrm{entropy}(p)<\textrm{entropy}(q)-\alpha\big{)}$ $(1-\delta)\cdot q(x)+\delta\cdot p(x)$ Speculative
SpecCascade [OPT01] $\bm{1}\big{(}\max_{v}q(v)<\max_{v}p(v)-\alpha\cdot D_{\textrm{TV}}(p,q)\big{)}$ $(1-\delta)\cdot q(x)+\delta\cdot p(x)$ Speculative
SpecCascade [OPTLog] $\bm{1}\big{(}\,\textrm{entropy}(p)<\textrm{entropy}(q)-\alpha\cdot D_{\textrm{TV}}(p,q)\big{)}$ $(1-\delta)\cdot q(x)+\delta\cdot p(x)$ Speculative
Table 3: Target distributions associated with different inference algorithms, where $\alpha$ is a free parameter and $\beta\geq 1-\alpha$ is a parameter dependent on $q,p$ and $\alpha$. The last column indicates whether the execution is sequential (Algorithm 2), via an oracle (Algorithm 3), or speculative (Algorithm 5) (Leviathan et al., 2023). The third row presents a variant of the BiLD algorithm of Kim et al. (2023), where $D(q,p)$ is a measure of discrepancy between $q$ and $p$; the original algorithm differs from (Leviathan et al., 2023) in the use of a deterministic speculative decoding procedure with a dynamic draft window (see §5).
@@ Appendix B Derivation of Chow’s rule
The minimizer of equation 22 is of the form:
$r^{*}(x_{<t})=1\leavevmode\nobreak\ \leavevmode\nobreak\ \iff\leavevmode\nobreak\ \leavevmode\nobreak\ \mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},q_{t})\right]\,>\,\alpha.$ (23)
If $\ell=\ell_{\text{0-1}}$, one may employ a plug-in estimator to equation 23 by replacing the expected 0-1 loss over $q_{t}$ with $1-\max_{v}q_{t}(v)$, giving us $\hat{r}_{\rm\tt Chow}(x_{<t})$ in equation 2. If $\ell=\ell_{\log}$, one may replace the expected log loss over $q_{t}$ with the entropy of $q_{t}$, giving us:
$\hat{r}_{\rm\tt ChowLog}(x_{<t})=1\leavevmode\nobreak\ \iff\leavevmode\nobreak\ \textrm{entropy}\big{(}q(\cdot|x_{<t})\big{)}>\alpha,$ (24)
@@ Appendix C Optimal Deferral: Additional Discussion
We provide additional discussion for the optimal deferral rules derived in §3 and §4.
@@ C.1 Optimal sequential deferral when $\ell=\ell_{\log}$
$r^{*}(x_{<t})=1\leavevmode\nobreak\ \leavevmode\nobreak\ \iff\leavevmode\nobreak\ \leavevmode\nobreak\ \mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},q_{t})\right]\,>\,\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},p_{t})\right]+\alpha\cdot D_{\textup{{TV}}}(p_{t},q_{t}).$
@@
$\textstyle\hat{r}_{\rm\tt DiffLog}(x_{<t})=1\leavevmode\nobreak\ \leavevmode\nobreak\ \iff\leavevmode\nobreak\ \leavevmode\nobreak\ \sum_{v}q_{t}(v)\cdot\log(q_{t}(v))\,<\,\sum_{v}p_{t}(v)\cdot\log(p_{t}(v))-\alpha.$ (25)
@@ C.2 Optimal speculative deferral when $\ell=\ell_{\log}$
$r^{*}(x_{<t})=1\leavevmode\nobreak\ \leavevmode\nobreak\ \iff\leavevmode\nobreak\ \leavevmode\nobreak\ \mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},q_{t})\right]\,>\,\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},p_{t})\right]+\alpha\cdot D_{\textup{{TV}}}(p_{t},q_{t}).$
@@
$\textstyle\hat{r}_{\rm\tt OPTLog}(x_{<t})=1\leavevmode\nobreak\ \iff\leavevmode\nobreak\ \sum_{v}q_{t}(v)\cdot\log(q_{t}(v))\,<\,\sum_{v}p_{t}(v)\cdot\log(p_{t}(v))-\alpha\cdot D_{\textup{{TV}}}(p_{t},q_{t}).$ (26)
@@
$L_{\rm spec}(r_{\rm\tt OPT};x_{<t})-\min_{r}L_{\rm spec}(r;x_{<t})\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \textstyle B_{q}\cdot\sum_{v\in\mathscr{V}}\big{|}\mathbb{P}(v|x_{<t})-q_{t}(v)\big{|}\,+\,B_{p}\cdot\sum_{v\in\mathscr{V}}\big{|}\mathbb{P}(v|x_{<t})-p_{t}(v)\big{|}.$
@@
Plugging these bounds into the equivalent of equation 18 in Lemma 5 for the log-loss completes the proof. ∎
@@ C.3 Optimal speculative deferral for greedy decoding
When $T\rightarrow 0$, running Algorithm 5 with $\tilde{r}_{\rm\tt OPT}$ as the deferral rule and $\tilde{q}_{t}$ as the drafter is equivalent to running it with $\hat{r}_{\rm\tt Diff}$ in equation 5 as the deferral rule and $\tilde{q}_{t}$ as the drafter.
@@
Note that under greedy inference, $\tilde{q}_{t}$ $\tilde{p}_{t}$ are one-hot encodings of $\operatorname*{arg\,max}_{v}q_{t}(v)$ and $\operatorname*{arg\,max}_{v}p_{t}(v)$ respectively. As a result,
$D_{\textrm{TV}}(\tilde{q}_{t},\tilde{p}_{t})=\bm{1}\left(\operatorname*{arg\,max}_{v}q_{t}(v)\neq\operatorname*{arg\,max}_{v}p_{t}(v)\right).$
@@
[…]
$\pi(v)=1-\delta_{\rm OPT}(q,p)=\bm{1}\left(\max_{v}q(v)\,\geq\,\max_{v}p(v)-\alpha\right)=1-\delta_{\rm\tt Diff}(q,p).$
@@
When $\operatorname*{arg\,max}_{v}q(v)=\operatorname*{arg\,max}_{v}p(v)$, $p_{\rm res}(v)=0$. When $\operatorname*{arg\,max}_{v}q(v)\neq\operatorname*{arg\,max}_{v}p(v)$,
@@
C.4 Equivalence between equation 7 and equation 8
Since the prefix $x_{<t}$ is fixed in equation 7, the constrained optimization we seek to solve is of essentially of the following form:
$\min_{r\in\{0,1\}}(1-r)\cdot c_{0}+r\cdot c_{1}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \text{s.t.}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ r\cdot c_{2}\leq B,$